OpenStudy (anonymous):

Can you please help me?? How many liters of a 70% acid solution must be mixed with a 25% acid solution to get a 405L of a 60% acid solution?

OpenStudy (amistre64):

so you mean to tell me, that posting the same question multiple times isnt getting itself answered?

OpenStudy (anonymous):

well no one is helping me and I don't know how to solve this problem

OpenStudy (amistre64):

i can help you, but this type of online etiquette comes across as .. well, rude. All your doing is pushing yourself to the front of the line by it. Just saying. Now, lets see what the question is asking for

OpenStudy (amistre64):

we have 2 amounts that need to be added togather to obtain a total amount of 405 right?

OpenStudy (anonymous):

Being rude wasn't my intention.. so im sorry

OpenStudy (amistre64):

i know, but just keep it in mind :) we can label these amounts anything we want; ill call them a + b = 405

OpenStudy (anonymous):

ok and yeah just those two amounts

OpenStudy (amistre64):

something that might not be so obvious i that percentages multiply .... to wit i mean: a(.25) + b(.70) = 405(.60)

OpenStudy (amistre64):

a + b = 405 a(.25) + b(.70) = 405(.60) now this is the setup, and they usually have you do this becasue they have taught you about solving a system of equations prior

OpenStudy (amistre64):

Do you have any notion as to what should be done with this?

OpenStudy (anonymous):

do you multiply the 405 and .60, then do you add both of the numbers and then divide? I'm honestly not sure

OpenStudy (amistre64):

that is a pretty good assumption. But a method i prefer for this is called elimination; since they only want one value we can modify either equation to .. well... eliminate the part that is useless to us

OpenStudy (amistre64):

when we multiply or divide an equation; the value doesnt change, only the way it looks. a + b = 405 ; we can multiply this one by -.25, and hopefully youll see why: -a(.25) - b(.25) = 405(-.25) a(.25) + b(.70) = 405(.60) --------------------------- and add the results 0 b(.70-.25) = 405(.60-.25)

OpenStudy (amistre64):

now all we have to do is solve for the amount that is left; that "b" value since it was at 70% b(.70-.25) = 405(.60-.25) b = 405(.60-.25) / (.70-.25) and what do you get when you simplify that?

OpenStudy (anonymous):

im gonna say 315 but I think that's very wrong.. aren't you supposed to subtract the .70 and the .25, then you multiply 405(.35). Last do you divide 141.75 by .45?

OpenStudy (anonymous):

I must sound like an idiot

OpenStudy (amistre64):

315 is correct

OpenStudy (anonymous):

hmm wow

OpenStudy (amistre64):

we can dbl chk it if you like :) 90(.25)+315(.70) = 405(.60) 90(.25)+315(.70) - 405(.60) = 0

OpenStudy (amistre64):

if that works out your good :)

OpenStudy (anonymous):

yeah that works out perfectly, thank you very much I do appreciate it

OpenStudy (amistre64):

youre welcome :) and good luck

OpenStudy (anonymous):

thank you!